OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 3x + y = 1 6x + 2y = −9 0 2 There is no x value as there is no solution to this system x can be any value as there are infinitely many solutions to this system

OpenStudy (anonymous):

3x + y = 1 ---> y = 1 - 3x 6x + 2y = −9 6x + 2 (1 - 3x) = -9 6x + 2 - 6x = -9 6x - 6x + 2 = -9 2 = -9. This is false. So there is no solution. There is no x value as there is no solution to this system.